OpenStudy (samanthagreer):

What is the value of y? 6(5 + 3) = 6 • 5 + 6 • y i think y=3

OpenStudy (zaynab123):

yeah its 3

OpenStudy (youngstudier):

I believe you are correct @samanthagreer ^-^

OpenStudy (albert0898):

Step 1. Distribute 6(5 + 3) = 6 • 5 + 6 • y You get 30 + 18 = 6 * 5 + 6 * y Step 2. Combine like terms 48 = 30 + 6y Step 3. Begin to isolate the variable. Like terms with like terms. 18 = 6y 3 = y

OpenStudy (samanthagreer):

ok thanks